OpenStudy (anonymous):

if 3a=2b=5c then 1/a : 1/b : 1/c = ?

OpenStudy (anonymous):

Alright so let's give "a" , a random number, shall we? let's use a=2 so this means that 3*2 = 2b so 6=2b hence meaning that b=3 2b=5c so 2*3 = 5c so c = 6/5 = 1.2 now that you have values for a, b, and c, just find the ratio :)